Hi,
Can someone please help me solve 4y=-28+7x
I need to find the y and x intercepts and graph.
I do not understand how to solve this.
Pamela G

Answer by nerdybill(7384)

To find the y-intercept, set x=0 and solve for y:
4y=-28+7x
4y=-28+7(0)
4y=-28
y = -7
therefore, the y-intercept is at (0, -7)
.
.
To find the x-intercept, set y=0 and solve for x:
4y=-28+7x
4(0)=-28+7x
0=-28+7x
28=7x
4 = x
therefore, the x-intercept is at (4, 0)